1.) Explain the difference between average velocity and instantaneous velocity. Is it possible for these two values to be different?

2.) Given a position-time graph, describe how you would determine the average velocity and instantaneous velocity,
1
Expert's answer
2019-12-24T14:33:51-0500

1.)

The instantaneous velocity at a given point can be defined as slope of the tangent of the curve drawn at a point while average velocity is equal to the slope of the secant line which intersects the function at the beginning and end of the interval.

average velocity = "\dfrac{X_f-X_i}{t_f-t_i}"

instantaneous velocity= "\dfrac{d(x(t))}{dt}"

These two velocities can same or different.


2.) In a position-time graph, average velocity can be determined by the formula "\dfrac{X_f-X_i}{t_f-t_i}" or by finding the slope of the secant line which intersects the function at the beginning and end of the interval

and instantaneous velocity can be determined by "\dfrac{d(x(t))}{dt}" at a particular point